Circular No. 7461

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

δ SCORPII

J. Fabregat, University of Valencia; and P. Reig, Foundation for Research and Technology Hellas and University of Crete, Heraklion, write: “This star is undergoing a major optical and Hα outburst. Spectra (resolution 0.3 nm, range 580–760 nm) taken with the Skinakas Observatory 1.3-m telescope reveal the Hα line in emission. Measured equivalent widths are −0.34 ± 0.02 nm on July 19.89 UT and −0.33 ± 0.02 nm (FWHM 0.69 ± 0.01 nm) on July 20.90. We made observations after S. Otero, Buenos Aires, reported a gradual brightening (visual mag estimates: June 30.03, 2.24; July 4.08, 2.23; 16.09, 2.06; 20.02, 2.04). In 1990, Cote and van Kerkwijk (1993, A.Ap. 274, 870) observed weak emission on the flanks of the Hα line core in absorption, and they proposed the classification of δ Sco as a Be star. All other observations in the literature present a normal absorption-line spectrum, while our observations constitute the first detection of emission lines in δ Sco, confirming its Be-star nature.”

COMET C/1999 S4 (LINEAR) H. Weaver, Johns Hopkins University, on behalf of the Hubble Space

Telescope (HST) ToO comet team, reports the following results: “HST images of the comet show a dramatic increase in activity on July 5, with the flux in a 0′′.15 square aperture increasing by a factor of ∼ 1.5 in just under 4 hr (from July 5.776 to 5.940 UT). During HST observations one day later (July 6.717–6.889), the activity levels were decreasing and were ∼ 3 times lower for the final observation, compared to the peak value from the previous day. The flux in the last HST image (on July 7.961) was ∼ 7 times lower than the peak value measured on July 5. On July 7, at least one ‘fragment’ is seen 0′′.85 (460 km, projected) from the nucleus in the tailward direction, and a sharp tailward spike of emission is observed, reminiscent of the morphology observed during the outburst in C/1996 B2 (Hyakutake) in late-March 1996. HST spectroscopic data taken on July 5 with STIS show evidence for emissions from CO, C, H, O, and possibly S. S_{2}, CS, and OH were detected during STIS observations on July 6, and OH, CS, NH, and possibly S_{2} were detected on July 7. Preliminary production rates are 5 × 10^{26} (July 5), 1.4 × 10^{24} (July 6), and 1.2 × 10^{29} (July 6) for CO, S_{2}, and H_{2}O, respectively, but the CO and S_{2} values could change by a factor of ∼ 2 or so, pending final analysis. Nevertheless, one firm conclusion is that CO is strongly depleted in C/1999 S4, relative to the observed abundances in C/1996 B2 and C/1995 O1.”

2000 July 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green